June 17-20

– Five stores provided ~62.5 hours of luminosity
– Cryo system technicians repaired the Tevatron sector F4 wet engine flywheel
– Pbar and RF personnel resolved air-flow cooling problem for Accumulator RF station
– Booster RF station BRF6 bypassed and replaced with BRF19
– Two transfers of antiprotons were lost
– Booster extraction kicker problem repaired
– NuMI beam tripped off when helium flow through the target increased – it was reset
– NuMI personnel conducted target scans on Monday, June 20. They will move the target today.

Read the Current Accelerator Update
Read the Early Bird Report
View the Tevatron Luminosity Charts